# GAPrompt: Geometry-Aware Point Cloud Prompt for 3D Vision Model

### v2

@@ 3 The Proposed Method
We illustrate our GAPrompt method for efficiently fine-tuning pre-trained 3D vision models in detail, which contains three geometry-aware prompt modules: the Point Prompt, the Point Shift Prompter, and the Prompt Propagation mechanism. As shown in Figure 3, given a pre-trained 3D transformer with $N$ blocks and a specific downstream task, we freeze the backbone and solely update newly introduced GAPrompt modules and the classification head.
@@ 3.1 Point Prompt
Given a raw input point cloud $\boldsymbol{x}\in\mathbb{R}^{S\times 3}$ with $S$ points, firstly we hybrid Point Prompt $\boldsymbol{\mathcal{P}}\in\mathbb{R}^{P\times 3}$ into its 3D space, denoted as $[\boldsymbol{x};\boldsymbol{\mathcal{P}}]\in\mathbb{R}^{(S+P)\times 3}$, where “[][\;][ ]” indicates concatenation and $P$ is the number of learnable points. Additionally, we modify the shape of $\boldsymbol{x}$ with our Point Shift Prompter, producing a shifted point cloud $\tilde{\boldsymbol{x}}\in\mathbb{R}^{S\times 3}$. This module also generates instance-specific informative shape features $\boldsymbol{f}\in\mathbb{R}^{D}$, where $D$ is the embedding dimension of transformers, formulated as:
@@
Then the hybrid point cloud $[\boldsymbol{x};\boldsymbol{\mathcal{P}}]$ becomes prompted input point cloud $[\tilde{\boldsymbol{x}};\boldsymbol{\mathcal{P}}]\in\mathbb{R}^{(S+P)\times 3}$.
Following the original architecture of the pre-trained model, the prompted point cloud is encoded into $L_{t}$ point tokens $\boldsymbol{h}_{1}$ by the token embedding module. We denote input tokens of $i$-th block as $\boldsymbol{h}_{i}\in\mathbb{R}^{L_{t}\times D}$. After that, input tokens $\boldsymbol{h}_{i}$ are concatenated with $L_{p}$ prompt tokens $\boldsymbol{p}_{i}\in\mathbb{R}^{L_{p}\times D}$ enhanced by shape feature $\boldsymbol{f}$, formulated as $[\boldsymbol{h}_{i};\boldsymbol{p}_{i}]$. Then we feed these tokens into our Prompt Propagation mechanism, injecting prompt tokens into the feature extraction process:
@@ 3.2 Point Shift Prompter
where $\tilde{\boldsymbol{d}}_{j}^{n}\in\mathbb{R}^{C_{j}\times K_{j}\times D_{j}}$ is features of neighbor points $n_{j}$ in $j$-th level. After upsampling from level $k$ to level $1$, the feature propagation process yields $\tilde{\boldsymbol{d}}_{1}^{n}$. We concatenate $\tilde{\boldsymbol{d}}_{1}^{n}$ and $\tilde{\boldsymbol{d}}_{1}$ and send into a small MLP, namely shift head, to generate the $\tilde{\boldsymbol{x}}$:
@@ 3.3 Prompt Propagation
Given a set of input tokens $[\boldsymbol{h}_{i};\boldsymbol{p}_{i}]\in\mathbb{R}^{(L_{t}+L_{p})\times D}$ in the $i$-th block, we can find the geometric relationship between input tokens via FPS and KNN. We formulate it as follows:
@@ B.1 Prompt Injection
Given a set of input tokens and prompt tokens $[\boldsymbol{h}_{i};\boldsymbol{p}_{i}]\in\mathbb{R}^{(L_{t}+L_{p})\times D}$ in the $i$-th block, we can find geometric relationship between input tokens via FPS and KNN algorithm. We use $\boldsymbol{h}_{i}^{c}\in\mathbb{R}^{C\times D}$ and $\boldsymbol{h}_{i}^{n}\in\mathbb{R}^{C\times K\times D}$ to denote center tokens and local neighboring tokens, where $c$ and $n$ represent the index for center and neighboring tokens. $L_{t}$ is the input tokens number, $L_{p}$ is the prompt tokens number, and $C$ and $K$ represent the number of local centers and local neighbors respectively.
@@ Appendix C Additional Experiments
C.1 Analysis on Point Prompt Number ${P}$ and Initialization.
We perform quantitative analysis on Point Prompt about two additional different initialization distributions and learnable point number ${P}$. We respectively test initializing all points in a uniform distribution and in cluster forms, where each cluster follows a Gaussian distribution in 3D space. And we explore the impact of different point numbers. As shown in Figure 8, The result turns out that uniform distribution generally outperforms the cluster distribution, and with 20 learnable points, the model achieves the best results. This suggests that appropriate additional point prompts help to better capture subtle geometric features and to acquire explicit awareness of geometry.

### v3

@@ 3 The Proposed Method
We illustrate our GAPrompt method for efficiently fine-tuning pre-trained 3D vision models in detail, which contains three geometry-aware prompt modules: the Point Prompt, the Point Shift Prompter, and the Prompt Propagation mechanism. As shown in Figure 3, given a pre-trained 3D transformer with NN blocks and a specific downstream task, we freeze the backbone and solely update newly introduced GAPrompt modules and the classification head.
@@ 3.1 Point Prompt
Given a raw input point cloud $\boldsymbol{x}\in\mathbb{R}^{S\times 3}$ with $S$ points, firstly we hybrid Point Prompt $\boldsymbol{\mathcal{P}}\in\mathbb{R}^{P\times 3}$ into its 3D space, denoted as $\boldsymbol{x};\boldsymbol{\mathcal{P}}]\in\mathbb{R}^{(S+P)\times 3}$, where “[][\;]” indicates concatenation and $P$ is the number of learnable points. Additionally, we modify the shape of $\boldsymbol{x}$ with our Point Shift Prompter, producing a shifted point cloud $\tilde{\boldsymbol{x}}\in\mathbb{R}^{S\times 3}$. This module also generates instance-specific informative shape features $\boldsymbol{f}\in\mathbb{R}^{D}$, where $D$ is the embedding dimension of transformers, formulated as:
@@
Then the hybrid point cloud $\boldsymbol{x};\boldsymbol{\mathcal{P}}]$ becomes prompted input point cloud $\tilde{\boldsymbol{x}};\boldsymbol{\mathcal{P}}]\in\mathbb{R}^{(S+P)\times 3}$.
Following the original architecture of the pre-trained model, the prompted point cloud is encoded into $L_{t}$ point tokens $\boldsymbol{h}_{1}$ by the token embedding module. We denote input tokens of $i$-th block as $\boldsymbol{h}_{i}\in\mathbb{R}^{L_{t}\times D}$. After that, input tokens $\boldsymbol{h}_{i}$ are concatenated with $L_{p}$ prompt tokens $\boldsymbol{p}_{i}\in\mathbb{R}^{L_{p}\times D}$ enhanced by shape feature $\boldsymbol{f}$, formulated as $\boldsymbol{h}_{i};\boldsymbol{p}_{i}]$. Then we feed these tokens into our Prompt Propagation mechanism, injecting prompt tokens into the feature extraction process:
@@ 3.2 Point Shift Prompter
where $\tilde{\boldsymbol{d}}_{j}^{n}\in\mathbb{R}^{C_{j}\times K_{j}\times D_{j}}$ is features of neighbor points $n_{j}$ in $j$-th level. After upsampling from level $k$ to level 11, the feature propagation process yields $\tilde{\boldsymbol{d}}_{1}^{n}$. We concatenate $\tilde{\boldsymbol{d}}_{1}^{n}$ and $\tilde{\boldsymbol{d}}_{1}$ and send into a small MLP, namely shift head, to generate the $\tilde{\boldsymbol{x}}$:
@@ 3.3 Prompt Propagation
Given a set of input tokens $\boldsymbol{h}_{i};\boldsymbol{p}_{i}]\in\mathbb{R}^{(L_{t}+L_{p})\times D}$ in the $i$-th block, we can find the geometric relationship between input tokens via FPS and KNN. We formulate it as follows:
@@ B.1 Prompt Injection
Given a set of input tokens and prompt tokens $\boldsymbol{h}_{i};\boldsymbol{p}_{i}]\in\mathbb{R}^{(L_{t}+L_{p})\times D}$ in the $i$-th block, we can find geometric relationship between input tokens via FPS and KNN algorithm. We use $\boldsymbol{h}_{i}^{c}\in\mathbb{R}^{C\times D}$ and $\boldsymbol{h}_{i}^{n}\in\mathbb{R}^{C\times K\times D}$ to denote center tokens and local neighboring tokens, where $c$ and $n$ represent the index for center and neighboring tokens. $L_{t}$ is the input tokens number, $L_{p}$ is the prompt tokens number, and $C$ and $K$ represent the number of local centers and local neighbors respectively.
@@ Appendix C Additional Experiments
C.1 Analysis on Point Prompt Number P{P} and Initialization.
We perform quantitative analysis on Point Prompt about two additional different initialization distributions and learnable point number P{P}. We respectively test initializing all points in a uniform distribution and in cluster forms, where each cluster follows a Gaussian distribution in 3D space. And we explore the impact of different point numbers. As shown in Figure 8, The result turns out that uniform distribution generally outperforms the cluster distribution, and with 20 learnable points, the model achieves the best results. This suggests that appropriate additional point prompts help to better capture subtle geometric features and to acquire explicit awareness of geometry.